### v2

# Review and Prospect of Algebraic Research in Equivalent Framework between Statistical Mechanics and Machine Learning Theory

Abstract
Mathematical equivalence between statistical mechanics and machine learning theory has been known since the 20th century, and researches based on such equivalence have provided novel methodology in both theoretical physics and statistical learning theory. For example, algebraic approach in statistical mechanics such as operator algebra enables us to analyze phase transition phenomena mathematically.
In this paper, for theoretical physicists who are interested in artificial intelligence, we review and prospect algebraic researches in machine learning theory. If a learning machine has hierarchical structure or latent variables, then the random Hamiltonian cannot be expressed by any quadratic perturbation because it has singularities. To study an equilibrium state defined by such a singular random Hamiltonian, algebraic approach is necessary to derive asymptotic form of the free energy and the generalization error. We also introduce the most recent advance, in fact, theoretical foundation for alignment of artificial intelligence is now being constructed based on algebraic learning theory.
This paper is devoted to the memory of Professor Huzihiro Araki who is a pioneer founder of algebraic research in both statistical mechanics and quantum field theory.
1 Introduction
Statistical mechanics was founded in the 19th century and have evolved throughout the 20th and 21th centuries. Information theory was established in the 20th century [59], and formal similarity between statistical mechanics and information theory has been sometimes pointed out [38, 17, 42, 13]. If nature and principle of two fields are different, such formal similarity can be treated by common mathematical foundation.
In the second half of the 20th century, equivalent structure was found between statistical mechanics of random interactions and learning and memory in artificial neural networks [20, 60, 54, 36, 47, 44, 62]. The extensive researches in these field became one of the foundations of today’s artificial intelligence.
Among mathematical physics in statistical mechanics and quantum field theory, algebraic approach was indicated in the 1960s by pioneer wrok by Huzihiro Araki [8, 9, 10]. In fact, types of von Neumann algebra made by a spin system was classified, quantum field theory can be treated by axiomatic way, and mathematical equivalence of equilibrium state and variational principle was clarified.
In this paper, for theoretical physicists who are interested in artificial intelligence, we review and prospect algebraic researches in machine learning theory and its relation to artificial intelligence alignment. We expect that machine learning theory would be extended from the viewpoint of statistical physics.
This paper consists of four sections. In the second section, we explain the well-known formal equivalence of statistical physics and machine learning theory. In subsection 2.1, structure of statistical physics is described with which physicists is familiar, however, this subsection is prepared for comparison with the following subsections. In subsection 2.2, structure of machine learning theory is introduced. Physicists understand that machine learning theory is equivalent to statistical mechanics with random Hamiltonian. In subsection 2.3, we explain how concepts in statistical mechanics correspond to ones in machine learning theory. In the third section, we review and prospect algebraic approach in machine learning theory. If a learning machine contains hierarchical structure or hidden variables, then the random Hamiltonian has singularities, resulting that algebraic approach is necessary. In subsection 3.1, the main results in algebraic approach of machine learning theory are summarized. It is shown that asymptotic form of the free energy and generalization error are represented by two birational invariants, the real log canonical threshold and singular fluctuation. In subsection 3.2, mathematical properties of real log canonical threshold are clarified and its application to estimation of the free energy is explained. In subsection 3.3, mathematical properties of singular fluctuation are studied and its application to estimation of the generalization error is examined. In subsection 3.4, the most recent researches about artificial intelligence alignment is prospected. We expect that mathematical researches about statistical physics and machine learning theory would be a foundation of the future artificial intelligence development.
2 Equivalent Framework of Statistical Mechanics and Statistical Learning Theory
In this section we summarize the well-known common framework of statistical mechanics and statistical learning theory [20, 60, 54, 36, 47, 44, 62]. Even if statistical mechanics and statistical learning theory are based on the different nature and principle, their formal structure can be studied based on the common mathematical framework. From the mathematical point of view, there is no need to distinguish them.
2.1 Statistical Mechanics of Random Hamiltonian
First, we explain the formal framework of statistical mechanics, which is well-known in statistical physicists, however, this subsection is necessary for comparing with the following subsections.
Let $J$ be a random variable defined on a finite dimensional Euclidean space and $w\in\mathbb{R}^{d}$ be an observable. In order to examine a physical system with random interaction, a function $H(w,J)$ is set as a candidate model, which is referred to as a random Hamiltonian. For a given random variable $J$, the equilibrium state with an inverse temperature $\beta>0$ is defined by a probability density function of $w$,
$\rho(w|J)=\frac{1}{Z(\beta,J)}\exp(-\beta H(w,J))\varphi(w),$ (1)
where $\varphi(w)$ is a nonnegative function of $w$ and $Z(\beta,J)$ is the normalizing constant,
[…]
If the principle of equal weights is adopted, then $\varphi(w)\equiv 1$. The normalizing constant $Z(\beta,J)$ is called a partition function. The expectation and variance with respect to the probability distribution $\rho(w|J)$ are denoted by $\mathbb{E}_{w}^{\beta}[\;\;]$ and $\mathbb{V}_{w}^{\beta}[\;\;]$, respectively. The free energy is defined by
$F(\beta,J)=-\frac{1}{\beta}\log Z(\beta,J).$
The average free energy over $J$ is denoted by $\mathbb{E}_{J}[F(\beta,J)]$. It is well known in statistical mechanics that, if the average free energy is explicitly calculated as a function of $\beta$, then several important values can be derived from the average free energy. Validation of modeling a random Hamiltonian is performed by comparing the theoretical results with the experimental observation.
Example in Statistical Mechanics. Let $J=\{J_{ij}\}$ and $w=\{w_{i}\}$. For example, as a probability distribution of $J$, a normal distribution is sometimes employed. A physical model of a spin glass system [20, 60] is defined by a random Hamiltonian,
$H(w,J)$ =$=-\sum_{\{i,j\}}J_{ij}w_{i}w_{j}-\sum_{i}h_{i}w_{i}$ (2)
where $h=\{h_{i}\}$ is the external magnetic field. If the principle of equal wights is adopted, $\varphi(w)\equiv 1$. Then the conditional distribution $\rho(w|J)$ represents Boltzmann distribution of the equilibrium state with inverse temperature $\beta>0$. If the average free energy is calculated, the average magnetization is obtained from the free energy,
$\chi(\beta)=\mathbb{E}_{J}\mathbb{E}_{w}^{\beta}\left[\frac{1}{n}\sum_{i}w_{i}\right]=\frac{1}{n}\sum_{i}\frac{\partial}{\partial h_{i}}\mathbb{E}_{J}[F_{n}(\beta,J)].$
The appropriateness of a model given by eq.(2) can be validated by comparing the average magnetization $\chi(\beta)$ with experimental results.
2.2 Statistical Learning Theory
Second, we explain machine learning theory. Statistical physicists can understand that machine learning theory is formally equivalent to statistical mechanics.
Let $d$, $n$, $N$ are positive finite integers. In stead of random interaction $J$, we use a notation for a random sample $X^{n}$. Let
$X^{n}\equiv\{X_{i}\;;\;i=1,2,...,n\}$
be a set of $\mathbb{R}^{N}$ valued random variables which are independently subject to a probability density function $q(x)$. A learning machine is defined by a conditional probability density function $p(x|w)$ and a probability density function $\varphi(w)$ where $x\in\mathbb{R}^{N},w\in\mathbb{R}^{d}$. In statistical learning theory, $q(x)$ is an unknown data-generating distribution, and a pair of $p(x|w)$ and $\varphi(w)$ is a candidate model which is designed by a user. There is no guarantee that a model will be appropriate for an unknown data-generating process. Therefore we need mathematical theory which holds for an arbitrary triple $(q(x),p(x|w),\varphi(w))$.
A random Hamiltonian
$H(w,X^{n})=-\sum_{i=1}^{n}\log p(X_{i}|w),$ (3)
is called the minus log likelihood in learning theory. The probability density function
$\rho(w|X^{n})=\frac{1}{Z(\beta,X^{n})}\varphi(w)\prod_{i=1}^{n}p(X_{i}|w)^{\beta}$
is called a posterior distribution, where
$Z(\beta,X^{n})=\int\varphi(w)\prod_{i=1}^{n}p(X_{i}|w)^{\beta}dw.$
This definition is equal to eq.(1) because $\exp(-\beta H(w,X^{n}))=\prod_{i=1}^{n}p(X_{i}|w)^{\beta}$. The expectation and variance using $\rho(w|X^{n})$ are also denoted by $\mathbb{E}_{w}^{\beta}[\;\;]$ and $\mathbb{V}_{w}^{\beta}[\;\;]$, respectively. The inference methods $\beta=1$ and $\beta=\infty$ are called Bayesian and maximum likelihood ones, respectively. The free energy is given by
$F(\beta,X^{n})=-\frac{1}{\beta}\log Z(\beta,X^{n}).$
Then random variables $Z(1,X^{n})$ and $F(1,X^{n})$ are referred to as the marginal likelihood and the minus log marginal likelihood, both of which are important values in statistics. Also $F(1,X^{n})$ is called stochastic complexity or Bayesian code length in information theory. Note that $Z(1,X^{n})$ is a probability density function
[…]
and the average free energy is equal to the entropy and the relative entropy.
[…]
where $q(x^{n})=\prod_{i=1}^{n}q(x_{i})$. This equation shows that the smaller free energy is equivalent to the smaller relative entropy. The posterior predictive distribution is defined by the average of $p(x|w)$ in the equilibrium state,
$p(x|X^{n})=\mathbb{E}_{w}^{\beta}[p(x|w)].$
Then the generalization error is defined by the relative entropy of $q(x)$ and $p(x|X^{n})$,
$G_{n}(\beta)=\int q(x)\log\frac{q(x)}{p(x|X^{n})}dx,$
which measures a difference between the data-generating function $q(x)$ and the posterior predictive distribution $p(x|X^{n})$. The smaller generalization error shows that a pair $p(x|w)$ and $\varphi(w)$ is more appropriate for the unknown data-generating distribution. Hence, one of the most important purposes in statistical learning theory is to clarify the behavior of the generalization error $G_{n}(\beta)$ as a function of $n$. However, the integration over $q(x)$ in the definition $G_{n}(\beta)$ cannot be performed, because $q(x)$ is unknown. On the other hand, the training error
$T_{n}(\beta)=\frac{1}{n}\sum_{i=1}^{n}\log\frac{q(X_{i})}{p(X_{i}|X^{n})}$
is used to estimate the generalization error. One of the purposes of theoretical study is to derive the difference between the generalization and training errors
$\mathbb{E}_{X^{n}}[G_{n}(\beta)-T_{n}(\beta)]$
[…]
which is not equal to zero in general because $X_{i}$ and $X^{n}$ are not independent. If $\beta=1$, then
$p(X_{n+1}|X^{n})=\frac{Z(1,X^{n+1})}{Z_{n}(1,X^{n})},$
resulting that
$\mathbb{E}_{X^{n}}[G_{n}(1)]$ $=\mathbb{E}_{X^{n+1}}[F(1,X^{n+1})-F(1,X^{n})]-S,$ (5)
where
[…]
is the entropy of the data-generating distribution. The relation eq.(5) shows that the generalization error is derived from the free energy if it is given by an explicit function of $n$. In machine learning, for the validation of candidate models $p(x|w)$ and $\varphi(w)$, $F(1,X^{n})$ and $G_{n}(1)$ are examined.
Example. Let $x=(x_{1},x_{2})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ and assume that $\{(X_{1i},X_{2i});i=1,2,...n\}$ is a set of independent random variables whose probability density function is $q(x_{1},x_{2})$. For the statistical learning for $q(x_{2}|x_{1})$
$p(x|w)=\frac{q(x_{1})}{(2\pi\sigma^{2})^{N/2}}\exp\left(-\frac{1}{2\sigma^{2}}(x_{2}-f(x_{1},w_{1}))^{2}\right)$ (6)
where $f(x,w_{1})$ is a function from $\mathbb{R}^{M}$ to $\mathbb{R}^{N}$. In deep learning, a function $f(x,w)$ is defined by using a layered neural network. The parameter is $w=(w_{1},\sigma)$. Remark that, in this model, $q(x_{1})$ is not estimated.
2.3 Correspondence between Physics and Machine Learning
In this subsection, we summarize the correspondence between statistical mechanics and machine learning, which is displayed in Table 1.
In statistical mechanics, the random Hamiltonian is devised as a candidate model for studying a phenomenon which has random interaction such as spin glass systems. Sometimes an assumption is set that $J$ is subject to normal distribution, validation of which is performed by comparing theoretical values with experimental observation. In statistical mechanics, it is sometimes said that, even if a candidate model is not completely constructed, the essential property of natural phenomenon can be captured by renormalizing procedure, which is referred to as the universality of theoretical physics.
In Bayesian statistics [25, 26, 46], a pair of $p(x|w)$ and $\varphi(w)$ is a candidate model for estimating unknown information source. In statistics, it is premised that all models are wrong [16, 15]. If a pair is designed, then $X^{n}$ is a set of exchangeable random variables. For exchangeable random variables, by de Finetti-Hewitt-Savage theorem [33], there exists a functional probability distribution $Q$ such that
$q(x)$ $\sim Q,$ (7)
$X^{n}$ $\sim\prod_{i=1}^{n}q(x_{i}).$ (8)
[…]
In statistical mechanics, thermodynamical limit plays the important roles, by which a macroscopic phenomenon is derived from a microscopic stochastic dynamics. For example, the dimension of $w=\{w_{i}\}$ is made to be infinity in spin glass theory. Also in machine learning theory, thermodynamical limit can taken if the leaning machine has homogeneous structure. In fact, if the function $f(x_{1},w_{1})$ in eq.(6) is given by the inner product of $x_{1}$ and $w_{1}$,
$f(x_{1},w_{1})=(x_{1})\cdot(w_{1})$
then it is possible to study thermodynamical limit. However, non-homogeneous physical phenomena such as Bose-Einstein condensation, it is impossible to study thermodynamical limit. In this paper we study nonhomogeneous learning machines such as hierarchical and degenerate neural networks which contain complex singularities, resulting that there may not exist the thermodynamical limit. It is the problem for the future study how to examine the large scale limit of nonhomogeneous learning machines. Almost all learning machines are singular [72] and that’s good [84], which may be one of the most essential properties of brain-like inference systems.
Statistical Mechanics Machine Learning Theory
$J$ random interaction $X^{n}$ random data
$w$ observable $w$ parameter of model
$H(w,J)$ random Hamiltonian $H(w,X^{n})$ minus log likelihood
$\varphi(w)$ weight of $w$ $\varphi(w)$ prior distribution of $w$
$\rho(w|J)$ equilibrium state $\rho(w|X^{n})$ posterior distribution
$Z(\beta,J)$ partition function $Z(\beta,X^{n})$ marginal likelihood
$F(\beta,J)$ free energy $F(\beta,X^{n})$ stochastic complexity
Table 1: Statistical Mechanics and Learning Theory
3 Algebraic Researches in Statistical Learning Theory
In this chapter we review and prospect algebraic researches in statistical learning theory.
3.1 Mathematical Results
In this subsection, we explain the main results of singular learning theory.
Assume that $\varphi(w)$ is a $C_{0}^{\infty}$ class function and that $W\subset\mathbb{R}^{d}$ is a compact subset whose open kernel $W^{o}$ contains the support of $\varphi(w)$, $\mbox{supp}\;\varphi$. Let us define a function of $W$
[…]
Then the expectation value of the random Hamiltonian eq.(3) is
$\mathbb{E}_{X^{n}}[H(w,X^{n})]=nL(w).$
Note that $L(w)\geq S$, where $S$ is the entropy of $q(x)$. We assume that $L(w)$ is a real analytic function of $w\in W^{o}$ and that there exists $w_{0}\in\mbox{supp}\;\varphi$ which minimizes $L(w)$ $(w\in W)$. The set of all parameters which minimize $L(w)$ in $W$ is denoted by
$W_{0}\equiv\{w\in W\;;\;L(w)=L(w_{0})\}.$
In other words, $W_{0}$ consists of all zero points of the real analytic function $L(w)-L(w_{0})$, which is called an analytic set. If $w_{0}$ satisfies $q(x)=p(x|w_{0})$, then $q(x)$ is said to be realizable by $p(x|w)$. Note that, if $q(x)$ is realizable by $p(x|w)$, then $L(w_{0})$ is equal to the entropy $S$ of $q(x)$. If there exists a unique $w_{0}$ that minimizes $L(w)$ and if the Hessian matrix $\nabla^{2}L(w_{0})$ is positive definite, then $q(x)$ is said to be regular for $p(x|w)$, if otherwise it is said to be singular for $p(x|w)$.
If $q(x)$ is regular for $p(x|w)$, then the posterior distribution concentrates in the neighborhood of $w_{0}$ as $n\rightarrow\infty$, resulting that the posterior distribution can be asymptotically approximated by a normal distribution,
$\rho(w|X^{n})\approx(1/C)\exp\left(-\frac{\beta n}{2}(w-w^{*})(\nabla^{2}L(w_{0}))(w-w^{*})\right),$
where $w^{*}$ is the unique parameter that minimizes $H(w,X^{n})$. The conventional theory for regular case is called regular learning theory. Statistical theory for singular case is called singular learning theory which contains regular theory as a special case. If a learning machine contains latent variables or hierarchical structure, then the regularity condition is not satisfied [30, 29, 22], hence we need singular learning theory both in statistics and machine learning. For example, in deep learning, the rank of the Hessian matrix in the neighborhood of the obtained parameter by training is far smaller the dimension of $w$.
Let $w_{0}$ be an element of the set $W_{0}$. We define a function
$K(w)=L(w)-L(w_{0}).$
We need a method how to analyze the set
$W_{\varepsilon}\equiv\{w\in W\;;\;K(w)\leq\varepsilon\}$
for a real analytic function $K(w)$. In general, the set $W_{0}$ contains singularities, hence we need singularity theory to study $W_{0}$. Such a method was constructed based on singularity theory,algebraic geometry, and algebraic analysis [11, 40] . The following theorem is the foundation of singular learning theory. Hironaka Resolution Theorem [34]. For a given real analytic function $K(w)\geq 0$, there exist both a compact subset ${\cal M}$ of a $d$-dimensional analytic manifold and a proper real analytic function from ${\cal M}$ to $W$
$g:{\cal M}\ni u\mapsto g(u)\in W$
such that, in each local coordinate of ${\cal M}$, $K(g(u))$ is normal crossing, in other words,
$K(g(u))$ $=u_{1}^{2k_{1}}u_{2}^{2k_{2}}\cdots u_{d}^{2k_{d}},$ (9)
$\varphi(g(u))|g^{\prime}(u)|$ $=b(u)|u_{1}^{h_{1}}u_{2}^{h_{2}}\cdots u_{d}^{h_{d}}|,$ (10)
where $k=(k_{1},k_{2},...,k_{d})$ and $h=(h_{1},h_{2},...,h_{d})$ are multi-indices of nonnegative integers, in which at least one $k_{i}$ is a positive integer. Here $b(u)>0$ is a positive analytic function, and $|g^{\prime}(u)|$ is the absolute value of the Jacobian determinant of $w=g(u)$. The correspondence between $W\setminus W_{0}$ and $g^{-1}(W\setminus W_{0})$ is one-to-one. Note that a function $w=g(u)$ is called proper if the inverse image of a compact set is also compact. The real log canonical threshold (RLCT) $\lambda$ and multiplicity $m$ are defined by
$\lambda$ $=\min_{L.C.}\min_{1\leq j\leq d}\left(\frac{h_{j}+1}{2k_{j}}\right),$ (11)
$m$ $=\max_{L.C.}\#\left\{j;\frac{h_{j}+1}{2k_{j}}=\lambda\right\},$ (12)
where $\min_{L.C.}$ and $\max_{L.C.}$ show the minimum and maximum values over all local coordinates respectively. Here we define $(h_{j}+1)/(2k_{j})=\infty$ for $k_{j}=0$, and ##\## means the number of elements of a set. Hence $0<\lambda<\infty$ and $1\leq m\leq d$. Since $W$ is compact and $w=g(u)$ is proper, the number of all local coordinates can be taken to be finite. The property and application of RLCT is discussed in subsection 3.2.
Additionally we assume the relatively finite condition that, there exists $c_{0}>0$ such that, for an arbitrary $w\in W$,
[…]
where $f(x,w)=\log(p(x|w_{0})/p(x|w))$. Since $W$ is a compact set, both sides of eq.(13) are trivially finite. Hence eq.(13) requires the inequality for the case $f(x,w)\rightarrow 0$, in other words, the variance of $f(x,w)$ is bounded by its average. In the case when this condition is not satisfied, the variance of Hamiltonian cannot be bounded by its average, resulting that the free energy has a different asymptotic behavior. [50]. By using resolution theorem and the relatively finite condition, it is proved that the following limit is finite
$\nu(\beta)=\frac{\beta}{2}\lim_{n\rightarrow\infty}n\;\mathbb{E}_{X^{n}}\left[\mathbb{V}_{w}^{\beta}[\log p(X_{i}|w)]\right],$ (14)
which is referred to as singular fluctuation. The property and application of singular fluctuation is discussed in subsection 3.3.
The main results of singular learning theory [73, 77] are described as follows.
$\mathbb{E}_{X^{n}}[F_{n}(\beta)]$ $=nK(w_{0})+nS+\frac{\lambda}{\beta}\log n-\frac{m-1}{\beta}\log\log n+O(1),$ (15)
$\mathbb{E}_{X^{n}}[G_{n}(\beta)]$ $=K(w_{0})+\frac{1}{n}\left(\frac{\lambda-\nu(\beta)}{\beta}+\nu(\beta)\right)+o\left(\frac{1}{n}\right),$ (16)
$\mathbb{E}_{X^{n}}[T_{n}(\beta)]$ $=K(w_{0})+\frac{1}{n}\left(\frac{\lambda-\nu(\beta)}{\beta}-\nu(\beta)\right)+o\left(\frac{1}{n}\right).$ (17)
These equations show that the free energy and generalization and training errors are asymptotically given by the real log canonical threshold and singular fluctuation. It follows that
$\mathbb{E}_{X^{n}}[G_{n}(\beta)]=\mathbb{E}_{X^{n}}\left[T_{n}(\beta)+\beta\mathbb{V}_{w}^{\beta}[\log p(X_{i}|w)]\;\right]+o\left(\frac{1}{n}\right)$ (18)
holds without explicit values of $\lambda$ and $\nu(\beta)$, which shows that the difference between the generalization error and the training error is asymptotically equal to the posterior variance of point-wise log likelihood function. Hence we can compare their theoretical values with the experimental ones.
3.2 Real log canonical threshold
In this subsection, mathematical properties and applications of the real log canonical threshold are studied. The concept log canonical threshold is a well-known birational invariant in high dimensional complex algebraic geometry [41]. The real log canonical threshold (RLCT) is the corresponding concept in real algebraic geometry [56, 45].
Let us introduce a zeta function of $z\in\mathbb{C}$,
[…]
which was firstly created by Gel’fand [23] and has been studied by many mathematicians. By the definition, $\zeta(z)$ is an holomorphic function in $\Re(z)>0$ which can be analytically continued to the unique meromorphic function on the entire complex plane. This unique analytic continuation can be proved by either the Hironaka resolution theorem in subsection 3.1 or the existence of Bernstein-Sato polynomial [14]. There exist both a differential operator $D_{w}$ and a polynomial $b(z)$ such that, for arbitrary $z\in\mathbb{C}$ and $w\in W$,
$D_{w}K(w)^{z+1}=b(z)K(w)^{z}.$
Then the monic polynomial that has the lowest order and satisfies this equation is called Bernsterin-Sato polynomial. By Hironaka resolution theorem, All poles of the zeta function is real and negative numbers. The largest pole and its order of the zeta function are equal to $(-\lambda)$ and $m$ respectively. By using the zeta function, it is trivial that statistical learning theory of a pair
$(p(x|w),\varphi(w))$
is equivalent to that of a pair
$(p(x|g(u)),\varphi(g(u))|g^{\prime}(u)|),$
hence both $(-\lambda)$ and $m$ are birational invariants.
It was also proved by Hironaka that there exists an algebraic algorithm by which both ${\cal M}$ and $w=g(u)$ can be found by finite recursive blow-ups [34]. In general, such ${\cal M}$ and $w=g(u)$ are not unique. If Newton diagram of $K(w)$ is nondegenerate, they can be found by using a toric modification, which was applied to statistics and machine learning [89].
RLCT is determined uniquely for a given pair $(K(w),\varphi(w))$. There are several mathematical properties.
  1. 1.
    If there exists the unique $w_{0}\in W_{0}$ such that $\nabla^{2}L(w_{0})$ is positive definite, and $\varphi(w_{0})>0$, then $\lambda=d/2$.
  2. 2.
    If there exists $w_{0}\in W_{0}$ such that $\det(\nabla^{2}L(w_{0}))=0$ and $\varphi(w_{0})>0$, then $20<\lambda<d/$ 2.
  3. 3.
    Note that Jeffreys’ prior is equal to zero at singularities, and if Jeffreys’ prior is employed in singular models, then $\lambda\geq d/2$.
  4. 4.
    Assume that $\lambda_{j}$ $(j=1,2)$ are RLCTs of $(K_{j}(w_{j}),\varphi_{j}(w_{j}))$. Then
    • RLCT of $(\sum_{j}K_{j}(w_{j}),\prod_{j}\varphi_{j}(w_{j}))$ is equal to $(\sum_{j}\lambda_{j})$.
    • RLCT of $(\prod_{j}K_{j}(w_{j}),\prod_{j}\varphi_{j}(w_{j}))$ is equal to $(\min_{j}\lambda_{j})$.
  5. 5.
    Assume that $\lambda_{j}$ $(j=1,2)$ are RLCTs of $(K_{j}(w),\varphi_{j}(w))$ and that $K_{1}(w)\leq c_{1}K_{2}(w)$ and $\varphi_{1}(w)\geq c_{2}\varphi_{2}(w)$ for some $c_{1},c_{2}>0$. Then $\lambda_{1}\leq\lambda_{2}$.
  6. 6.
    Let $K_{1}(w)=\sum_{j=1}^{J_{1}}f_{j}(w)^{2}$ and $K_{2}(w)=\sum_{j=1}^{J_{2}}g_{j}(w)^{2}$. If the ideal generated from $\{f_{j}(w)\}$ is equal to that from $\{g_{j}(w)\}$, then $(K_{1}(w),\varphi_{1}(w))$ and $(K_{2}(w),\varphi_{2}(w))$ have the same RLCT.
By using these properties, RLCTs of important statistical models and learning machines were found by developing resolution procedures in neural networks [69, 7] deep linear networks [6], deep convolution ReLU network [51] with and without skip connection, normal mixtures [85], Poisson mixtures [57], multinomial mixtures [83], general and nonnegative matrix facorizations [5, 31], Boltzmann machines [87], hidden and general Markov models [86, 92] and latent Dirichlet allocations [32].
[…]
If $q(x)$ is regular for $p(x|w)$, then $\lambda=d/2$ and $m=1$, hence $F(1,X^{n})$ is approximated by BIC [58],
${\rm BIC}=H(\hat{w},X^{n})+\frac{d}{2}\log n,$
where $\hat{w}$ is the parameter which minimizes $H(w,X^{n})$. In regular cases, the difference between the free energy and BIC is a constant order random variable. For general cases when $q(x)$ may be singular for for $p(x|w)$, the singular BIC was proposed [19] by using the estimated RLCT $\hat{\lambda}$,
${\rm sBIC}=H(\hat{w},X^{n})+\hat{\lambda}\log n.$
The difference between $F(1,X^{n})$ and sBIC is smaller than any $\log n$ order random variable.
Another method for approximation of the free energy was proposed [76]. Since $F(0,X^{n})=0$, there exists $10<\beta^{*}<$ 1 such that
$F(1,X^{n})$ $=\frac{\partial F}{\partial\beta}(\beta^{*},X^{n})=\mathbb{E}_{w}^{\beta^{*}}[H(w,X^{n})].$ (19)
Then we can show that
$\beta^{*}=1/\log n+o_{p}(1/\log n).$
By the definition,
${\rm WBIC}=\mathbb{E}_{w}^{\beta^{*}}[H(w,X^{n})],$
the difference between $F(1,X^{n})$ and $WBIC$ is smaller than any $\log n$ order random variable. In the numerical calculation of WBIC, the posterior distribution with the inverse temperature $1/\log n$ is necessary. An efficient algorithm to generate such posterior distribution in mixtures models were proposed [79].
3.3 Singular Fluctuation
In this subsection, mathematical properties and applications of singular fluctuation are studied. Singular fluctuation appears in many statistical contexts [1, 2, 3, 48].
The original appearance of the singular fluctuation is model selection criteria. if $q(x)$ is regular for $p(x|w)$, the singular fluctuation is equal to
$\nu(\beta)=\frac{1}{2}\mbox{tr}(IJ^{-1}),$
which does not depend on $\beta$, where $I$ and $J$ are $d\times d$ matrices,
[…]
$J$ $=-\nabla^{2}L(w_{0}).$ (21)
Moreover, if $q(x)$ is realizable by $p(x|w)$, then $I=J$, resulting that $\nu(\beta)=d/2$ where $d$ is the dimension of $w$.
If $q(x)$ is regular for $p(x|w)$, the limit $\beta\rightarrow\infty$ results in the maximum likelihood method in statistics. The posterior distribution converges to the delta function on the maximum likelihood estimator. Then
$\mathbb{E}_{X^{n}}[G_{n}(\infty)]=\mathbb{E}_{X^{n}}[T_{n}(\infty)]+d/n$
holds, which was firstly found by Akaike [1]. The $2n$ times of the right hand side of this equation is called Akaike information criterion. If $q(x)$ is not realizable by $p(x|w)$, then $d$ is replaced by $\mbox{tr}(IJ^{-1})$ which is called by Takeuchi information criterion. For the case when $\beta=1$ and $q(x)$ is regular for and realizable by a model, the $\mathbb{E}_{X^{n}}[G(1,X^{n})]$ can be estimated by deviance information criterion (DIC) [61]. The right hand side of eq.(18) is called WAIC which can be used to estimate the generalization error in general cases.
The leave-one-out cross validation (LOOCV) [24, 63, 27] is an alternative method to estimate the generalization error.
$C_{n}(\beta)=-\frac{1}{n}\sum_{i=1}^{n}\log p(X_{i}|X^{n}\setminus X_{i}),$
where
$X^{n}\setminus X_{i}=\{X_{j}\;;\;j=1,2,...,n\;(j\neq i)\}$
is the sample leaving $X_{i}$ out. Then the generalization error can be estimated by $C_{n}(\beta)$,
$\mathbb{E}_{X^{n}}[C_{n}(\beta)]=\mathbb{E}_{X^{n-1}}[G_{n-1}(\beta)]+S$
Note that $C_{n}(\beta)$ can be numerically calculated by using the posterior average,
$C_{n}(\beta)=-\frac{1}{n}\sum_{i=1}^{n}\log\left(\frac{\mathbb{E}_{w}^{\beta}[p(X_{i}|w)^{1-\beta}]}{\mathbb{E}_{w}^{\beta}[p(X_{i}|w)^{-\beta}]}\right).$ (22)
The difference between LOOCV and WAIC is smaller than any $1/n$ order random variable. In numerical calculation of eq.(22), the stability of the expectation $\mathbb{E}_{w}^{\beta}[p(X_{i}|w)^{-\beta}]$ is not ensured if a leverage sample point is contained in a sample [55, 21]. A new calculation method of LOOCV was proposed by using the approximation in MCMC distribution [64]. When the conditional probability is estimated in regression problems, LOOCV requires that the sample is independent, whereas information criteria do only conditional independence. On the other hand, information criteria needs that the sample size $n$ is sufficiently large, hence LOOCV and information criteria are different statistical tools [77, 80]. The concept of generalization error was extended onto the problems such as covariate shift, causal inference [37], and overparametrized cases [53], and the generalized information criteria are now being developed.
3.4 Singular Learning Theory and AI Alignment
In 1966, a statistician I. J. Good, who firstly found the importance of marginal likelihood $Z(1,X^{n})$ in statistical inference, speculated as [28]
“An ultra-intelligent machine is a machine that can far surpass all the intellectual activities of any man however clever. The design of machines is one of these intellectual activities; therefore, an ultra-intelligent machine could design even better machines.”
Artificial neural network research started in the 1960s. Nowadays, artificial intelligence (AI) research is progressing so rapidly that it is difficult to predict to which direction AI takes us. It is said that alignment of artificial intelligence is necessary for safety and welfare of human being in the near future, where AI alignment is sometimes defined as AI working without deviating from the designer’s intention.
In order to consider AI alignment, we need all opinions from all different viewpoints, and problems to be studied are beging proposed [4, 12]. Note that singularities in learning machines make the free energy and the generalization loss smaller if Bayesian inferences, hence degenerate property of hierarchical learning machines can be considered to be an advantage [84]. In AI alignment, we need to study developmental landscape of in-context learning in also transfomers [35]. Deneracy in the loss function [43] is useful in mechanistic interpretability [18].
In this subsection, we discuss AI alignment from the viewpoint of equivalence of statistical mechanics and machine learning. There are at least three problems caused by singularities in artificial neural networks: lack of identifiability, difficulty in design of the prior distribution, and phase transition in learning process. These three points originated from the hierarchical structure of learning machines. Remark that hierarchical structure is necessary for a learning machine to have an ability of universal approximation, that is to say, an arbitrary continuous function can be approximated by multi-layered neural networks with higher precision. Moreover it makes the generalization error far smaller if a learning machine is over-parametrized. Therefore, the following three problems should be solved.
The first problem is lack of identifiability. A statistical model or a leaning machine $p(x|w)$ is called identifiable if the map $w\mapsto p(x|w)$ is one-to-one. Many of classical statistical models are identifiable, hence the role of the parameter is uniquely determined. Therefore, by checking the obtained parameter, we can examine the reason why a machine answers an output for a given inputs. However, in singular learning machines, the role of the parameter is not uniquely determined, hence even by checking the parameter, we cannot examine whether the model’s learning is subject to the designer’s intention.
By introducing an equivalence relation
$w_{1}\sim w_{2}\Longleftrightarrow(\forall x)\;\;p(x|w_{1})=p(x|w_{2}),$
and preparing the equivalence class $W/\sim$, the map from a parameter to probability density function can be made one-to-one. This method is called blow-down in algebraic geometry. However, if $w_{0}$ is a singularity in the original parameter space, then its neighborhood in $W/\sim$ has infinite dimension, resulting that difficulty in understanding the parameter cannot be solved by $W/\sim$.
The second problem is difficulty in design of prior distribution. In older Bayesianism in the 20th century, it was premised that uncertainty could be captured by a statistical model and the personal belief should be represented by a prior distribution. However, because of nonidentiability, each element of the parameter does not have any concrete meaning, hence the prior distribution cannot be determined by any a priori personal knowledge about data-generating process. One might think that Jeffreys’ prior may be chosen as an objective one, however, it makes the generalization error far larger than others, which destroys the advantage of hierarchical structure.
We need modern Bayesian statistics, in which a statistical model or a learning machine is only a candidate and that a prior distribution is not any belief of a person but a part of a statistical model. In the practical applications, ridge or lasso type prior distribution with hyperparameter is sometimes employed and the hyperparameter is optimized by several criteria including cross validation.
The third problem is the phase transition phenomenon in learning process. In classical statistical models, the optimal parameter $w_{0}$ is unique and the posterior distribution can be approximated as a normal distribution which converges to the delta function $\delta(w-w_{0})$ as the sample size increases. The posterior distribution may be understood as a gradual increase of confidence.
On the other hand, in deep learning, even if the sample size is huge, it is far smaller than the infinity for the deep neural network. Let $q(x)$ be unrealizable by $p(x|w)$. By using division of unity, there is a set of nonngegative $C_{0}^{\infty}$ class functions $\{\varphi_{k}(w)\}$ such that
$\varphi(w)=\sum_{k=1}^{K}\varphi_{k}(w)$
Then it follows that
$Z_{n}(\beta)=\sum_{k=1}^{K}Z_{nk}(\beta),$
where
[…]
Let $w_{k}$ be the minimum point of $L(w)$ in the support of $\varphi_{k}(w)$ and $\lambda_{k}$ be the real log canonical threshold for $(q(x),p(x|w),\varphi_{k}(w))$. The local free energy is given by
$F_{nk}(\beta)==nL(w_{k})+\frac{\lambda_{k}}{\beta}\log n+O(\log\log n)$
and
$F_{n}(\beta)=\min_{k}\left\{nL(w_{k})+\frac{\lambda_{k}}{\beta}\log n\right\}+O(\log\log n).$
Therefore the free energy is given by the minimum of the local free energy. In statistical learning theory, $L(w_{k})$ and $\lambda_{k}\log n$ are called bias and variance, respectively. Note that $\exp(-F_{nk})$ is in proportion to the posterior probability of the local coordinate. Therefore the optimal local parameter set for minimum free energy is automatically chosen by the posterior probability. However, it is not equal to the optimal parameter set for the minimum generalization error. The local parameter set chosen by the posterior distribution jumps from a neighborhood to another one as sample size invreases. Such jumping process makes difficult to diagnose whether the learning process is out of order or not.
Phase transition phenomena can be also observed in the different situation. In mixture models, a Dirichlet distribution is often chosen for a prior distribution of the mixture ratio. Then the posterior distribution has a phase transition according to the hyperparameter of Dirichlet distribution [77, 83]. In the different phases, RCLTs are different and the supports of asymptotic posterior distributions are different [77].
4 Conclusion
In this paper, for theoretical physicists who are interested in artificial intelligence, we introduced mathematical equivalence between statistical mechanics of random Hamiltonian and machine learning theory. Professor Huzihiro Araki indicated that algebraic concepts such as operator algebras play important role in statistical mechanics and quantum field theory. In this paper, we proposed that, also in machine learning theory, algebraic research is useful in both fundamental and practical fields. For example, such researches are becoming more important in developing statistics and artificial intelligence.
Acknowledgment
The author would like to express his sincere gratitude to Professor Huzihiro Araki who is a pioneer founder of algebraic research in both statistical mechanics and quantum field theory.